OpenStudy (anonymous):

=2. =-3, =5 ,||u||=1,||v||=2,||w||=7 find

OpenStudy (anonymous):

\[v=<2\cos\theta,2\sin\theta>\] any more information given?

OpenStudy (anonymous):

yes, just pick a part of question which i don't know how to do, I am editing the question

OpenStudy (anonymous):

\[<u+v,v+w>=<u+v,v>+<u+v,w>\\ \quad=<u,v>+<v,v>+<u,w><v,w>\]

OpenStudy (anonymous):

the second part is what i ask on previous post. I stuck there only

OpenStudy (anonymous):

<v.v? = ||v|| = 2 so?

OpenStudy (anonymous):

yep. \[<u+v,v+w>=(2)+(2)+(5)+(-3)=6\]

OpenStudy (anonymous):

not that, ||v||= \[\sqrt{<v,v>}\] =2

OpenStudy (anonymous):

---> <v,v>=4

OpenStudy (anonymous):

aah